Use the following scenario to answer the following questions:

Natasha can produce either 5,000 pounds of cheese or 20 houses per year. Jameson can produce either 5,000 pounds of cheese or 10 houses per year.

If this economy produces more than 20 houses per year, the opportunity cost of a house will ________ because ________.

a increase; an input with a higher opportunity cost must be used
b increase; all gains from trade have already been exploited
c remain the same; Natasha and Jameson produce in constant ratios regardless of level of output
d decrease; production will be specialized
e decrease; inputs will be obtained at lower cost

Countries engaged in international trade to specialize in production based on

a relative exchange rates.
b comparative advantage.
c relative inflation rates.
d relative levels of GDP.
e relative unemployment rates.
Answer #1

Answer 1 : If this economy produce more than 20 houses per year than the opportunity cost of the house will increase because an input with higher opportunity cost has been increased.​​​​​​

As if Natasha produced than the opportunity cost of producing one house = 5000/20= 250 pounds of cheese

Where as if Jameson produced house than opportunity cost of producing 1 house = 5000/10= 500 pounds of cheese

Answer2 : Option B is correct. Countries engaged in international trade to specialised in production based on comparative advantage. Comparative advantage means making comparison in the production of the goods. The ability of an economy to produce the good more efficiently than the other economy is known as comparative advantage.